The length of a rectangle is 5 inches more than twice the width. If the area is to be at least 33 square inches, what are the possibilities for the width?
Answer by Boreal(15235)

width=x
length=x+5
x(2x+5)>=33
2^2+5x-33>=0
(2x+11)(x-3)>=0
only positive root is at least x=3 inches width and (2x+5)=11 inches length
the width has to be at least 3 inches.